# Learning for Dynamic Combinatorial Optimization without Training Data

### v1

@@ C.1 DyCO-GNN with SAGEConv for dynamic Maxcut and MIS
C.2 Results of the last checkpoints of DyCO-GNN on dynamic TSP
@@
Method 500/0.38 1000/0.76 2000/1.51 3000/2.27 5000/3.78 10000/7.57 20000/15.16
st70 PI-GNN (static) 2.10292 2.11124 2.03993 1.94988 1.70642 1.55780 1.46919
PI-GNN (warm) 1.48778 1.49261 1.47614 1.46951 1.44118 1.43268 1.43030
DyCO-GNN (emb) 1.49505 1.49011 1.44635 1.45302 1.41491 1.38462 1.35803
DyCO-GNN (GNN) 1.65917 1.58705 1.46435 1.43776 1.44108 1.35839 1.30869
DyCO-GNN (full) 1.67305 1.68751 1.50680 1.46486 1.39795 1.37431 1.33007
C.3 Snapshot-level ApRs bar plots
@@
Appendix D Proof and extension of Theorem 1
D.1 The Goemans-Williamson (GW) algorithm
Let $G=(V,E)$ be an unweighted graph with $|{V}|=n$ nodes, and let $L\in\mathbb{R}^{n\times n}$ be its Laplacian matrix. Let $c^{*}$ denote the maximum achievable cut size on this graph. Our goal is to attain this maximum cut size.
@@ D.1 The Goemans-Williamson (GW) algorithm
Goemans and Williamson [13] showed that the best approximation ratio (of at least 0.87856 times the optimal value) for the MaxCut problem can be attained by proceeding as follows. We first reformulate the above QUBO problem as the Semidefinite Program (SDP):
@@
In more detail, consider the eigen-decomposition $X^{*}_{\text{SDP}}=A\Lambda A^{T}$ where $A\in\mathbb{R}^{n\times n}$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix, with the eigenvalues of $X^{*}_{\text{SDP}}$ on its diagonal. Define the matrix $Y=A\Lambda^{\frac{1}{2}}$; this can be viewed as an “embedding” of the solution produced by the SDP. In it, each row $Y_{i}$ corresponds to node $i$, which needs to be mapped to $+1$ or $-1$ (this will be the label of the node, which in turn will determines the two sets produced by the cut). To attain these node labels, in the rounding cut, the GW rounding algorithm samples a random vector $r\sim\mathcal{N}(0,I_{n})$ (where $I_{n}$ denotes the $n\times n$ identity matrix), and defines the cut to be $x_{i}=\text{sign}(Y_{i}^{T}r)$.
D.2 Proof of Theorem 1
@@
Next, we show that the projection $\text{Proj}_{\mathcal{X}}$ is continuous. This is because the space $\mathbb{S}^{n}$ is a finite-dimensional real Hilbert space under the Frobenius inner product $<A,B>=\text{Tr}(AB)$. The feasible set $\mathcal{X}\subset\mathbb{S}^{n}$ is the intersection of the cone of positive semidefinite matrices and an affine subspace defined by linear constraints $\text{diag}(X)=1$. Both the positive semidefinite cone and the affine subspace are closed and convex, and therefore $\mathcal{X}$ is a closed, convex subset of a Hilbert space. By the Hilbert Projection Theorem [5, Theorem 3.12] the Euclidean projection onto any closed convex subset is continuous. Therefore, the projection function $\text{Proj}_{\mathcal{X}}$ is continuous.
@@ D.3 Extension
The proof is straightforward: given the assumption that the SDP solver $\Pi(\cdot)$ is locally continuous, that means that small perturbations of the initial SDP solution $X_{0}$ map to perturbations of the solution $\Pi(X_{0})$. Consequently, the results of Theorem 1 are applicable. We note that SDP solvers for solving MaxCut, such as interior-point methods, are locally continuous under mild assumptions; e.g. [43].
Appendix E Broader impacts
The DyCO-GNN framework introduces an efficient, UL approach for solving DCO problems without requiring offline training or labeled datasets. DyCO-GNN’s ability to efficiently adapt to evolving problem instance snapshots without optimization from scratch or external supervision enables more rapid and resource-efficient optimization in time-sensitive environments such as supply chain management, transportation systems, disaster response, and real-time network routing.
The potential societal benefits of DyCO-GNN are significant. By avoiding heavy training pipelines, DyCO-GNN can operate in low-resource settings and adapt on-the-fly to changes in problem structure, promoting computational sustainability. Furthermore, the ability to dynamically update solutions has implications for reducing operational delays and improving responsiveness in real-world systems.
However, the framework may also pose ethical and societal risks. There is the risk of dual-use. Also, the objective functions optimized by DyCO-GNN may not capture critical social or ethical constraints, such as fairness or safety, which could lead to unintended consequences if applied to human-centered systems. To mitigate these risks, we encourage future work on integrating human-in-the-loop constraints and ethical guardrails directly into dynamic optimization objectives.

### v2

@@ D.1 DyCO-GNN with SAGEConv for dynamic MaxCut and MIS
D.2 Results of the last checkpoints of DyCO-GNN on dynamic TSP
@@
Method 500/0.39 1000/0.77 2000/1.53 3000/2.29 5000/3.82 10000/7.66 20000/15.31
st70 PI-GNN (static) 1.25686 1.25882 1.26492 1.29193 1.25541 1.21451 1.18396
PI-GNN (warm) 1.19896 1.18582 1.20358 1.21865 1.19866 1.19128 1.21490
DyCO-GNN (emb) 1.19227 1.18662 1.16549 1.15118 1.17218 1.15700 1.15481
DyCO-GNN (GNN) 1.24525 1.21279 1.16314 1.16665 1.18421 1.15773 1.15031
DyCO-GNN (full) 1.21969 1.23395 1.16589 1.16746 1.15945 1.16328 1.14535
D.3 Results of Non-neural baselines
Here we include the results of non-neural approaches built into NetworkX (Hagberg et al., 2008). Each entry of Tables 8, 9, and 10 is of the form ApR (wall clock time). For dynamic MIS, we consider the method in Boppana and Halldórsson (2006); the implementation in NetworkX only works for graphs with fewer than several hundred nodes without a recursion error, so we only report the result on Infectious.
Our method works the best except for dynamic MIS on Infectious and dynamic TSP on st70. We note that the method in Boppana and Halldórsson (2006) is specifically designed for solving MIS, and the implementation in NetworkX cannot handle graphs with more than several hundred nodes, whereas our method and the baselines have broader applicability and can easily handle graphs with more than 10,000 nodes. It has also been shown in Schuetz et al. (2022) that PI-GNN performs on par or better than Boppana and Halldórsson (2006) on random d-regular graphs. For TSP on st70, the result of the greedy approach is better than that of Gurobi, which could be an edge case. Nevertheless, our method improves upon the PI-GNN baselines by a large margin.
Table 8: Comparison with non-neural baseline on dynamic MaxCut. Values closer to 1 are better ($\uparrow$).
Dataset Cut-based greedy PI-GNN (static) PI-GNN (warm) DyCO-GNN
Infectious 0.96921 (5.6s) 0.98888 (3.32s) 0.97186 (0.56s) 0.98947 (0.56s)
UC Social 0.98083 (933.16s) 0.99823 (3.53s) 0.99519 (1.18s) 0.99825 (0.24s)
DBLP Fail to find a solution in 10 hours 0.99095 (8.44s) 0.99039 (8.44s) 0.99319 (0.14s)
Table 9: Comparison with non-neural baseline on dynamic MIS. Values closer to 1 are better ($\uparrow$).
Dataset Boppana and Halldórsson (2006) PI-GNN (static) PI-GNN (warm) DyCO-GNN
Infectious 0.93825(6.88s) 0.87651 (3.33s) 0.50845 (0.56s) 0.88254 (1.12s)
Table 10: Comparison with non-neural baseline on dynamic TSP. Values closer to 1 are better ($\downarrow$).
Dataset Cost-based greedy PI-GNN (static) PI-GNN (warm) DyCO-GNN
burma14 1.19812 (<0.01s) 1.03358 (6.21s) 1.13860 (6.21s) 1.01811 (6.21s)
ulysses22 1.24475 (<0.01s) 1.08867 (6.37s) 1.17228 (0.64s) 1.08516 (6.37s)
st70 0.94963 (<0.01s) 1.07768 (27.84s) 1.11914 (27.84s) 1.04617 (27.84s)
D.4 Sensitivity analysis on SP parameters.
From the results reported in Tables 11 and 12, $0.2\leq\lambda_{\text{shrink}}\leq 0.4$ and $\lambda_{\text{perturb}}=0.1$ are good default choices. When $\lambda_{\text{shrink}}$ approaches 1, the ApRs will potentially be the highest early on, as this setting is close to a naive warm start; the ApRs may drop in this case when the number of epochs increases because the parameters closer to convergence on the previous snapshot without proper SP are not suitable for adaptation. We note that one should not confuse the performance with respect to epochs with the case in regular deep learning training: we are showing the number of epochs each snapshot is optimized for before adapting the parameters for the next snapshot, and the ApRs are computed across all snapshots. In general, a larger $\lambda_{\text{shrink}}$ gives better performance when the time budget is extremely tight, while a smaller $\lambda_{\text{shrink}}$ gives better performance when the time budget constraint is slightly relaxed. $\lambda_{\text{perturb}}$ has a smaller effect than $\lambda_{\text{shrink}}$. $\lambda_{\text{perturb}}=1$ is usually too aggressive; values like 0.1 and 0.01 typically give the best results.
Table 11: Mean ApR on dynamic MaxCut using UC Social achieved by DyCO-GNN with different SP parameters. Values closer to 1 are better ($\uparrow$).
Time budget (epochs/seconds)
$\lambda_{\text{shrink}}$ $\lambda_{\text{perturb}}$ 50/0.06 100/0.12 200/0.23 300/0.34 500/0.56 1000/1.12 2000/2.22 3000/3.33
0.2 0.1 0.99142 0.99747 0.99850 0.99862 0.99865 0.99872 0.99873 0.99875
0.4 0.1 0.99403 0.99724 0.99825 0.99841 0.99844 0.99848 0.99846 0.99848
0.6 0.1 0.99544 0.99805 0.99832 0.99830 0.99829 0.99807 0.99793 0.99782
0.8 0.1 0.99702 0.99738 0.99760 0.99744 0.99748 0.99742 0.99740 0.99740
0.4 1.0 0.98437 0.99459 0.99731 0.99777 0.99793 0.99793 0.99789 0.99785
0.4 0.1 0.99403 0.99724 0.99825 0.99841 0.99844 0.99848 0.99846 0.99848
0.4 0.01 0.99408 0.99732 0.99826 0.99836 0.99840 0.99847 0.99848 0.99846
0.4 0.001 0.99416 0.99732 0.99826 0.99835 0.99839 0.99845 0.99848 0.99845
Table 12: Mean ApR on dynamic MIS using UC Social achieved by DyCO-GNN with different SP parameters. Values closer to 1 are better ($\uparrow$).
Time budget (epochs/seconds)
$\lambda_{\text{shrink}}$ $\lambda_{\text{perturb}}$ 50/0.06 100/0.12 200/0.23 300/0.34 500/0.56 1000/1.12 2000/2.22 3000/3.33
0.2 0.1 0.48314 0.53766 0.68657 0.80957 0.88256 0.90916 0.91371 0.91607
0.4 0.1 0.54819 0.66505 0.80108 0.87436 0.91599 0.92047 0.92046 0.91987
0.6 0.1 0.70601 0.83711 0.90015 0.89537 0.88777 0.87972 0.87283 0.86969
0.8 0.1 0.84145 0.82497 0.80988 0.80543 0.80072 0.79774 0.79606 0.79350
0.4 1.0 0.59099 0.69377 0.81221 0.86738 0.90229 0.90993 0.90641 0.90533
0.4 0.1 0.54819 0.66505 0.80108 0.87436 0.91599 0.92047 0.92046 0.91987
0.4 0.01 0.54720 0.66718 0.80084 0.87241 0.91504 0.92001 0.91994 0.91921
0.4 0.001 0.54733 0.66713 0.80096 0.87235 0.91491 0.92020 0.92046 0.91994
D.5 Snapshot-level ApRs bar plots
@@
Appendix E Proof and extension of Theorem 1
E.1 The Goemans-Williamson (GW) algorithm
Let $G=(V,E)$ be an unweighted graph with |V|=n|{V}|=n nodes, and let $L\in\mathbb{R}^{n\times n}$ be its Laplacian matrix. Let $c^{*}$ denote the maximum achievable cut size on this graph. Our goal is to attain this maximum cut size.
@@ E.1 The Goemans-Williamson (GW) algorithm
Goemans and Williamson (1995) showed that the best approximation ratio (of at least 0.87856 times the optimal value) for the MaxCut problem can be attained by proceeding as follows. We first reformulate the above QUBO problem as the Semidefinite Program (SDP):
@@
In more detail, consider the eigen-decomposition $X^{*}_{\text{SDP}}=A\Lambda A^{T}$ where $A\in\mathbb{R}^{n\times n}$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix, with the eigenvalues of $X^{*}_{\text{SDP}}$ on its diagonal. Define the matrix $Y=A\Lambda^{\frac{1}{2}}$; this can be viewed as an “embedding” of the solution produced by the SDP. In it, each row $Y_{i}$ corresponds to node $i$, which needs to be mapped to +1+1 or 1-1 (this will be the label of the node, which in turn will determines the two sets produced by the cut). To attain these node labels, in the rounding cut, the GW rounding algorithm samples a random vector $r\sim\mathcal{N}(0,I_{n})$ (where $I_{n}$ denotes the $n\times n$ identity matrix), and defines the cut to be $x_{i}=\text{sign}(Y_{i}^{T}r)$.
E.2 Proof of Theorem 1
@@
Next, we show that the projection $\text{Proj}_{\mathcal{X}}$ is continuous. This is because the space $\mathbb{S}^{n}$ is a finite-dimensional real Hilbert space under the Frobenius inner product $A,B>=\text{Tr}(AB)$. The feasible set $\mathcal{X}\subset\mathbb{S}^{n}$ is the intersection of the cone of positive semidefinite matrices and an affine subspace defined by linear constraints $\text{diag}(X)=1$. Both the positive semidefinite cone and the affine subspace are closed and convex, and therefore $\mathcal{X}$ is a closed, convex subset of a Hilbert space. By the Hilbert Projection Theorem (Bauschke and Combettes, 2017, Theorem 3.12) the Euclidean projection onto any closed convex subset is continuous. Therefore, the projection function $\text{Proj}_{\mathcal{X}}$ is continuous.
@@ E.3 Extension
The proof is straightforward: given the assumption that the SDP solver $\Pi(\cdot)$ is locally continuous, that means that small perturbations of the initial SDP solution $X_{0}$ map to perturbations of the solution $\Pi(X_{0})$. Consequently, the results of Theorem 1 are applicable. We note that SDP solvers for solving MaxCut, such as interior-point methods, are locally continuous under mild assumptions; e.g., Shapiro (1988).